OpenStudy (anonymous):

x is not equal to -3; 6x to the second power minus 54 over 5x plus 15 help explain and solve i need to know how it's done

OpenStudy (anonymous):

You cannot really solve this. All you can do is re-write it in a different form.

OpenStudy (anonymous):

\[6x^2-\frac{54}{5x}+15\]?

OpenStudy (anonymous):

\[6x^2 - 54 \over 5x + 15\] Is my guess and you have to do it by long division.

OpenStudy (anonymous):

Sorry i meant simply

OpenStudy (anonymous):

it gives me an answer of 6x-3 over 5 but i am unsure how we got there

OpenStudy (anonymous):

which one of us has the correct 'starting problem' ?

OpenStudy (anonymous):

your do

OpenStudy (anonymous):

polpak

OpenStudy (anonymous):

Ok, so to simplify mine you must use long division.

OpenStudy (anonymous):

First lets get rid of a factor of 5. 5x + 15 = 5(x + 3) So we will leave a factor of 5 on the bottom. and just divide by the other factor. 6x -18 _____________ x + 3 ) 6x^2 + 0x - 54 -6x^2 -18x ---------- -18x - 54 +18x +54 ---------- 0 So we have that \[{6x^2 - 54 \over 5x + 15} = \frac{1}{5}\cdot(6x-18) = \frac{6x-18}{5}\]